# Energy dynamics in a class of local random matrix Hamiltonians

### v2

@@ Acknowledgements.
The authors are thankful for fruitful discussions with John Chalker, Weitao Chen, Anatoly Dymarsky, Joseph Eix, Rebecca Flint, Siddarth Jindal, Aleksei Khindanov, Tomaž Prosen, Rustem Sharipov, Roland Speicher, Anastasiia Tiutiakina, Iris Ulčakar, and Marko Žnidarič. We thank Karen TerHorst for assistance with the graphics. K.P. and T.I. acknowledge support from the National Science Foundation under Grant No. DMR-2143635. J.K. acknowledges support from the U.S. Department of Energy under Grant No. DE-SC0023692. N.P. acknowledges support from the Natural Sciences and Engineering Research Council of Canada. J.R. acknowledges financial support from the Royal Society through the University Research Fellowship No. 201101. We also acknowledge the hospitality of the Boulder Summer School 2023: “Non-Equilibrium Quantum Dynamics” with support from the National Science Foundation Grant No. DMR-1560837, where this project was conceived.
@@ Appendix A Counting on the Bethe lattice
In this section, we detail the counting of the number of words $aibi$ with $a\in[z]^{n}$ and $b\in[z]^{m}$, for some fixed $i$, that are fully reducible. To count these, we require that $ibi$ reduces to some irreducible word, say $c$, (equivalently site on the Bethe lattice) and $a$ reduces to the inverse of that word, $c^{-1}$, i.e. $c$ read backwards. Then, we sum over all points $c$. As an equation, we have
@@
Here, $\text{red}(\cdot)$ means apply the algebraic relations $i^{2}=\emptyset$ to all symbols in the word until no further reduction is possible. The function $\delta(\text{red}(a)-c^{-1})$ is $1$ if $a$ reduces to $c^{-1}$ and 00 otherwise, and the sum over $c$ runs over all fully irreducible words (equivalently sites on the Bethe lattice).
@@ Appendix B Numerical free convolution
Appendix C Argument for a Gaussian density of states
In this section we give a simple informal argument for the full large $L$ chain to have a Gaussian density of states for any independent local terms satisfying $\braket{h_{i}}=0$ and $\braket{h_{i}^{2}}=1$  [^4] We thank John Chalker for pointing out essentially the argument that follows to us.. The moments
@@ Appendix C Argument for a Gaussian density of states
which also tends to $L^{n/2}$. The same argument would apply for $|i_{k}-i_{l}|\geq\xi$ when $\xi\ll L$. The reason this argument is heuristic is that we need to first take $L\rightarrow\infty$ and then take $n\rightarrow\infty$. The number of pairings is $(n-1)!!$!, and so
$|P_{n}|\rightarrow L^{n/2}(n-1)!!.$! . (93)

### v3

@@ Acknowledgements.
The authors are thankful for fruitful discussions with John Chalker, Weitao Chen, Anatoly Dymarsky, Joseph Eix, Rebecca Flint, Siddharth Jindal, Alexey Khudorozhkov, Aleksei Khindanov, Tomaž Prosen, Rustem Sharipov, Roland Speicher, Anastasiia Tiutiakina, Iris Ulčakar, and Marko Žnidarič. We thank Karen TerHorst for assistance with the graphics. K.P. and T.I. acknowledge support from the National Science Foundation under Grant No. DMR-2143635. J.K. acknowledges support from the U.S. Department of Energy under Grant No. DE-SC0023692. N.P. acknowledges support from the Natural Sciences and Engineering Research Council of Canada. J.R. acknowledges financial support from the Royal Society through the University Research Fellowship No. 201101. We also acknowledge the hospitality of the Boulder Summer School 2023: “Non-Equilibrium Quantum Dynamics” with support from the National Science Foundation Grants No. DMR-1560837 and DMR-2328793, where this project was conceived.
@@ Appendix A Counting on the Bethe lattice
In this section, we detail the counting of the number of words aibiaibi with $a\in[z]^{n}$ and $b\in[z]^{m}$, for some fixed $i$, that are fully reducible. To count these, we require that ibiibi reduces to some irreducible word, say $c$, (equivalently site on the Bethe lattice) and $a$ reduces to the inverse of that word, $c^{-1}$, i.e. $c$ read backwards. Then, we sum over all points $c$. As an equation, we have
@@
Here, $\text{red}(\cdot)$ means apply the algebraic relations $i^{2}=\emptyset$ to all symbols in the word until no further reduction is possible. The function $\delta(\text{red}(a)-c^{-1})$ is $1$ if $a$ reduces to $c^{-1}$ and 0 otherwise, and the sum over $c$ runs over all fully irreducible words (equivalently sites on the Bethe lattice).
@@ Appendix B Numerical free convolution
Appendix C Argument for Gaussian moments
In this section we give a simple informal argument for the full large $L$ chain to have moments matching that of the Gaussian distribution for any independent local terms satisfying $\braket{h_{i}}=0$ and $\braket{h_{i}^{2}}=1$  [^4] We thank John Chalker for pointing out the beginnings of this argument to us.. The moments
@@ Appendix C Argument for Gaussian moments
which also tends to $L^{n/2}$. The same argument would apply for $|i_{k}-i_{l}|\geq\xi$ when $\xi\ll L$. The reason this argument is heuristic is that we need to first take $L\rightarrow\infty$ and then take $n\rightarrow\infty$. The number of pairings is (n1)!!(n-1)!!, and so
$|P_{n}|\rightarrow L^{n/2}(n-1)!!.$ (93)
Figure 10: Numerical evidence that only reducible words contribute to the moments as $q\rightarrow\infty$. Panels (a) and (b) are the two RMT models considered, and panels (c) and (d) are the respective fractional error between the number $R_{n}$ of reducible words of length $n$ and the RMT average moments $\braket{H^{n}}_{100}$, the latter being an empirical average over $100$ realizations. Darker blue is larger $q$; the fractional error systematically decreases with $q$.
@@
Appendix D Irreducible correlators vanish.
In this appendix, we give some numerical evidence that only irreducible words of the local terms survive the large $q$ limit. Consider the local terms $\{h_{j}\}_{j=1}^{L}$ with PBC so we have that $h_{i}h_{j}=h_{j}h_{i}$ for |ij|>1|i-j|>1 (regarding $1\equiv L$). We numerically compute the first few average moments, $\braket{H^{n}}$, of $L=4$ and $L=6$ chains in PBC and compare them against the number reducible words of length $n$. Since the RMT average moments of $H$ are a sum of all possible words, agreement of the two calculations is good evidence that the irreducible words vanish and only the reducible ones contribute to the moments as $q\rightarrow\infty$.
The number of reducible words, $R_{n}$, of $L$ terms in PBC having length $n$ are counted symbolically by generating all possible words, then performing two sorting and reduction steps repeatedly until the word can no longer be changed. We assume the word contains an even number of each term, since it will otherwise be trivially irreducible. We then sort the terms (letters) by moving smaller indices to the left, with the constraint that we can only swap adjacent terms $h_{i},h_{j}$ if $i\neq j\pm 1$. The sorted word is then simplified using $h_{j}^{2}=1$ by removing any pairs of adjacent identical terms. We then apply the same algorithm, but this time sorting the string largest to smallest, and again we remove paired adjacent terms. This is repeated until no further reductions are possible. If this results in an empty string we have found a reducible word. Otherwise we conclude the word is irreducible.
Fig. 10 panels (c) [corresponding to model (a)] and (d) [corresponding to model (b)] show the fractional error
$|\braket{H^{n}}_{100}-R_{n}|/R_{n}$ (95)
where $\braket{H^{n}}_{100}$ is an empirical average over $100$ realizations of the RMT. We see that the fractional error is on the order of one percent, and that the error systematically decreases with $q$.